# Developing a Thailand solar irradiance map using Himawari-8 satellite imageries and deep learning models

### v1

@@ 6 Acknowledgment
The authors thank Parinthorn Manomaisaowapak for developing the GHI estimation model (version-2) and NDR solution (Thailand) Co., Ltd. for implementing the solar map platform as well as Tanan Boonyasirikul for the initial experiments on the DL models. We appreciate the discussions with the Division of Solar Energy Development, Department of Alternative Energy Development and Efficiency (DeDe), Ministry of Energy, which provided supplementary solar data and a historical yearly solar map. This work was co-funded by the Chula Engineering Research grant 2023 and partially supported by Grants for Development of New Faculty Staff, Ratchadaphiseksomphot Endowment Fund.

### v2

@@ 6 Acknowledgment
The authors thank Parinthorn Manomaisaowapak for developing the GHI estimation model (version-2), Justin Jiraratsakul and NDR solution (Thailand) Co., Ltd. for implementing the solar map platform as well as Tanan Boonyasirikul for the initial experiments on the DL models. We appreciate the discussions with the Division of Solar Energy Development, Department of Alternative Energy Development and Efficiency (DeDe), Ministry of Energy, which provided supplementary solar data and a historical yearly solar map. This work was co-funded by the Chula Engineering Research grant 2023 and partially supported by Grants for Development of New Faculty Staff, Ratchadaphiseksomphot Endowment Fund.